how to find the angle in a quadrilateral that has 4 angles with the numbers 85, 69, (6x+22), and (x+37) please help 6th grader
Answer by stanbon(75887)

how to find the angle in a quadrilateral that has 4 angles with the numbers 85, 69, (6x+22), and (x+37) please help 6th grader
----
The sum of the four angles is 360 degrees
----
Equation:
85 + 69 + 6x+22 + x+37 = 360 degrees
---
213 + 7x = 360
7x = 147
x = 21
-----
Angle: 6x+22 = 6*21+22 == 126+22 = 148 degrees
Angle: x+37 = 21+37 = 58 degrees
===================================
Cheers,
Stan H.
==============================